An uncharged conducted spherical shells placed in the electric field of positive charge (q). What will be the net flux through the shell. What is the SI unit of electric flux?


1
Expert's answer
2021-06-07T09:31:54-0400

Electric Flux through a surface is defined as the surface integral of the electric field lines passing normally through the surface.

"\phi=\intop E.ds"


According to gauss's law, total electric flux through a closed surface equals the net charge enclosed in the surface divided by the permittivitty.

Consider a point charge q. Its net electric flux through any closed surface enlcosing it will be be equal to

"\phi =\frac{q}{\epsilon_0}"

This is independent of the choice of the shape and size of the closed surface.

The SI unit of electric flux is Volt - metres (V m)


For example



For charged spherical shell, all charge lies on the outer circumference of the surface.

Consider a charged spherical shell with outer radius b and inner radius a. The charge resides at a radius b from the centre. Consider a gaussian surface of radius a<r<b. Charge enclosed in this guassian surface will be zero.

Using gauss's law,

"\intop E.ds=\frac{q}{\epsilon_0}"

enclosed


E×2πrdr=0

Where r lies

"a<r<b"

E=0